OpenStudy (anonymous):

Caleb makes 70% of his free throws and is shooting free throws after practice. What is the probability that a) He makes the first free throw he shoots. b) He does not make a free throw until his 3rd attempt. c) He does not make a free throw until his 10th attempt. d) How many free throws should Caleb expect to shoot until he makes his first shot?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Frostbite @Hero

OpenStudy (anonymous):

thats not the asnwer,i have to answer each part

jimthompson5910 (jim_thompson5910):

this sounds like you'll be using the geometric distribution do you remember how to use this distribution?

OpenStudy (anonymous):

hi r u still here @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes do you remember how to use the geometric distribution

OpenStudy (anonymous):

hmm... idts

jimthompson5910 (jim_thompson5910):

you would use the formula Pr(X = k) = (1-p)^(k-1)*p

jimthompson5910 (jim_thompson5910):

in each and every case for this problem, the value of p stays the same p = 0.7

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how would i do this for a?

jimthompson5910 (jim_thompson5910):

for part a), k = 1 for part b) k = 3 for part c) k = 10 part d) ... we'll get to this later

jimthompson5910 (jim_thompson5910):

Pr(X = k) = (1-p)^(k-1)*p Pr(X = k) = (1-0.7)^(k-1)*0.7 Pr(X = 1) = (1-0.7)^(1-1)*0.7 Pr(X = 1) = (0.3)^(0)*0.7 Pr(X = 1) = 1*0.7 Pr(X = 1) = 0.7

jimthompson5910 (jim_thompson5910):

So this means that the chances of him making a free throw on the first attempt is 0.7 or 70% this makes sense because it's given that on any random trial, he makes 70% of his free throws

OpenStudy (anonymous):

Ok!

jimthompson5910 (jim_thompson5910):

repeat this but instead of k = 1, use k = 3 for part b

OpenStudy (anonymous):

So a is: (a) Pr(X = k) = (1-p)^(k-1)*p Pr(X = k) = (1-0.7)^(k-1)*0.7 Pr(X = 1) = (0.3)^(0)*0.7 Pr(X = 1) = 1*0.7 Pr(X = 1) = 0.7 <--

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

ok, so b

jimthompson5910 (jim_thompson5910):

now you're using k = 3

OpenStudy (anonymous):

can u write the formula again here?

jimthompson5910 (jim_thompson5910):

Pr(X = k) = (1-p)^(k-1)*p Pr(X = k) = (1-0.7)^(k-1)*0.7 Pr(X = 3) = (1-0.7)^(3-1)*0.7 Pr(X = 3) = ???

OpenStudy (anonymous):

umm

OpenStudy (anonymous):

its .3^(1,4)

OpenStudy (anonymous):

0.1853??

OpenStudy (anonymous):

wait no its .063!?

jimthompson5910 (jim_thompson5910):

yep, it's 0.063

OpenStudy (anonymous):

ok cool! c?

jimthompson5910 (jim_thompson5910):

so he has a 6.3% chance of making his first shot on the third attempt

jimthompson5910 (jim_thompson5910):

for part c, you use k = 10

jimthompson5910 (jim_thompson5910):

Pr(X = k) = (1-p)^(k-1)*p Pr(X = k) = (1-0.7)^(k-1)*0.7 Pr(X = 10) = (1-0.7)^(10-1)*0.7 Pr(X = 10) = ???

OpenStudy (anonymous):

Pr(X=10) = (0.3)^ 0.63 right

jimthompson5910 (jim_thompson5910):

no that doesn't look right, one sec

OpenStudy (anonymous):

kkk

jimthompson5910 (jim_thompson5910):

Pr(X = k) = (1-p)^(k-1)*p Pr(X = k) = (1-0.7)^(k-1)*0.7 Pr(X = 10) = (1-0.7)^(10-1)*0.7 Pr(X = 10) = (0.3)^(9)*0.7 Pr(X = 10) = 0.000019683*0.7 Pr(X = 10) = 0.0000137781 So the probability that he makes his first shot on the 10th try is 0.0000137781

OpenStudy (anonymous):

ok cool

jimthompson5910 (jim_thompson5910):

for the last one, you just need to compute 1/p

OpenStudy (anonymous):

how do i do that?

jimthompson5910 (jim_thompson5910):

p = 0.7 so 1/p = 1/0.7 = ??

OpenStudy (anonymous):

1.43

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

now round up to get 2

jimthompson5910 (jim_thompson5910):

so he should expect to make 2 shots when the first one will go in (on average)

OpenStudy (anonymous):

kk! thanks can u help me with one more prob?

OpenStudy (anonymous):

A quarterback for the Seattle Seahawks completes 54% of his passes. Let the random variable X be the number of passes completed in 20 attempts. a) What is the probability that the quarterback throws three passes before he gets his first completion (on his fourth attempt)? b) How many passes should the quarterback expect to make before he gets his first completion? c) What is the probability that he has to throw more than 5 passes to get his first completion? (a) (b) (c) p= 54 percent = 0.54

jimthompson5910 (jim_thompson5910):

so this sounds like another geometric distribution problem

jimthompson5910 (jim_thompson5910):

any ideas?